the age of two sisters are 11 years and 19 years. in how many years' time will the sum of their ages be 70 years. form an equation.

Answer by TimothyLamb(4379)

x = number of years in future
---
11 + x + 19 + x = 70
2x = 70 - 11 - 19
x = (70 - 11 - 19)/2
x = 20
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
Send and receive free text messages online:
https://sooeet.com